1/ The Following data was obtained while observing a reaction between nitric oxide and hydrogen:

F2(g) + 2CLO2(g) -----> 2FCLO2(g)

Experiment Initial(F2) Initial(CLO2) Initial rate

(mol/L) (mol/L) (mol/L*s)

1/ 0.10 0.010 1.2 x 10^-3

2/ 0.10 0.040 4.8 x 10^-3

3/ 0.20 0.010 2.4 x 10^-3

A/ write the rate law for this reaction

B/ calculate the rate constant for the reaction

C/ what is the order of this reaction

2/ consider the reaction N2(g) + 3H2(g) ----> 2NH3(g). At a particular moment during the reaction molecular hydrogen is reacting at the rate of 0.075 mole/liter*second. At what rate is ammonia (NH3) being formed? At what rate is molecular nitrogen reacting?

3/ Ethyl iodide decompose at a certain temperature in the gas phase as follows:

C2H5I(g) -----> C2H4(g) + HI(g)

A plot of in C2H5I versus time yields a straight line with a slope of -1.36x10^2 mol/L*min

A/ what is the order of this reaction

B/ write the integrated rate law for this reaction.

C/ Determine the rate constant for this reaction.

D/ Determine the half life for this reactions.

E/ Determine the concentration of ethyl iodide after 75 minutes if the initial concentration of ethyl iode was 0.25M
Answer #1

1/ The Following data was obtained while observing a reaction between nitric oxide and hydrogen:

F2(g) + 2CLO2(g) -----> 2FCLO2(g)

Experiment      Initial(F2)        Initial(CLO2)       Initial rate

                       (mol/L)             (mol/L)                 (mol/L*s)

1                      0.10                 0.010                   1.2 x 10^-3

2                      0.10                0.040                      4.8 x 10^-3

3                      0.20                0.010                       2.4 x 10^-3

Rate law for this reaction is :

Rate = k [F2]m [ClO2]n

k is rate constant, m and n are the orders with respect to F2 and ClO2

we find m and n

Rate exp 3/ Rate 1 = 2

2 = [F2]3m [ClO2]3n/ [F2]1m [ClO2]1n     ( k is cancelled from both sides since the experiment is same. )

2 = (0.20/0.10)m   ….(conc of ClO2 is same)

2 = 2m

Ln 2 = m ln 2

m = 1

order with respect to F2 = 1

lets find order with respect to ClO2

Rate (exp 2) / rate (exp 1) = 4

4 = (0.040/0.01)n

4 = 4n

ln 4 = n ln 4

n = 1

order with respect to ClO2 = 1

Rate law:

R = k [F2][ClO2]

We use exp 1 to calculate rate constant

1.2E-3 = k ( 0.10 ) ( 0.010)

k = 1.2

C. ) Overall order = m + n = 1+1 = 2

Question 2 )

Rate of decomposition of H2 = 0.075 mol/L.s

We know

-d[H2]/ 3 dt = d [NH3 ]/ 2 dt

(0.075 mol per L per s / 3)*2 = d [NH3 ]/ dt

d[NH3]/dt = 0.05 mol per L per s

Q. 3 . )

A.) For this reaction plot of ln [C2H5I] vs time is straight line that suggests the order of this reaction if first

B.) We write integrated rate law for first order

ln [A]t = -kt + ln [A]0

C) Rate constant

ln [A]t = -kt + ln [A]0

y              m x + c

m = slope = -k

k = - slope

its slope is given

k is rate constant

k = - (-1.36E2)

= 1.36 E2 min-1

Rate constant

k = 1.36 E2 min-1

D) half life :

For first order reaction

t ½ = 0.693 / k

= 0.693 / 1.36E2

= 5.1 E-3 min

E.)
We use integrated rate law:

ln [CH3CH2I]= - 1.36 E2 min-1 * 75 min + ln [0.25]

ln [CH3CH2I] = -10201.4

lets take antilog of both side

[CH3CH2I]= 0

So there will no any ethyl iodide remains after 75 minute